proove that vector A=3i+j-2k B=-i+3j+4k C=4i-2j-6k can form the sides of a triangle and find the length of the median of the triangle.
1
Expert's answer
2019-10-14T09:14:33-0400

Solution:

For these vectors, the following equality

A-B=C

Really, 3-(-1)=4, 1-3=-2, -2-4=-6. So, this vectors is the sides of a triangle.

We find the length of the median as a half-sum of vectors A and B

m=0.5(3-1)i+0.5(1+3)j+0.5(-2+4)k

m=i+2j+k

Ansver: m=i+2j+k